OpenStudy (anonymous):

Let f(x) = x – 6 and g(x) = x2 – x + 4. Find g(f(x)). x2 – 13x + 46 x2 – 11x + 34 x2 – 2x – 2 x2 – x – 2

OpenStudy (anonymous):

g(f(x))=(x-6)2-(x-6)+4=x2-12x+36-x+6+4=x2-13x+46 A

OpenStudy (anonymous):